ZyCo uses the straight-line method. Assets purchased between the 1st and 15th of the month are depreciated for the entire month; assets purchased after the 15th of the month are treated as though they were acquired the following month. On May 2, 20X1, ZyCo purchases furniture for $35,000 that it expects to last for 10 years; ZyCo expects the furniture to have a residual value of $3,000. What is the book value at the end of 20X3?

$25,400

$25,667

$22,400

$26,467

$24,500

$23,467
Answer #1

Depreciable value = 35000 - 3000 = 32000

No of months till 2003 end = 8 + 12 + 12 = 32 months out of a toal of 10 years or 120 months

Depreciation = 32000 * 32 / 120 = 8533

Book value at end of 2003 = 35000 - 8533 = 26467